# On the Source Model Key Agreement Problem

### v2

@@ 1 Introduction
  • We conjecture that $S(X;Y\|Z)\neq I(X;Y)$ in the XOR case (see Section 4). Showing this would require developing a new upper bound for the source model problem. Partial results are also provided in our attempt to develop a new upper bound.
The rest of this paper is organized as follows: Section 2 evaluates the upper bound $\Psi(X;Y\|Z)$ for some settings. Finally, Section 4 discusses the tightness of the upper bound in the XOR case. Additionally, in Appendix A, we compute the secret key capacity for a class of sources where $Z$ is a function of $X$ and $P_{Y|X}$ corresponds to an erasure channel.
Notation and Definitions: Random variables are represented by uppercase letters, whereas their realizations are indicated by lowercase letters. It is assumed that all random variables possess finite alphabets. Sets are denoted by calligraphic letters. The notation $X-\!\!\!\!\minuso\!\!\!\!-Y-\!\!\!\!\minuso\!\!\!\!-Z$ signifies that $X$ and $Z$ are conditionally independent given $Y$, which implies that $P_{XZ|Y}=P_{X|Y}P_{Z|Y}$. In this scenario, we assert that $X-\!\!\!\!\minuso\!\!\!\!-Y-\!\!\!\!\minuso\!\!\!\!-Z$ constitutes a Markov chain. The sequences $(U_{1},U_{2},\ldots,U_{i})$ and $(U_{i},U_{i+1},\cdots,U_{n})$ are denoted as $U^{i}$ and $U_{i}^{n}$ respectively. For a sequence $X^{n}=(X_{1},X_{2},\cdots,X_{n})$, we use $X_{\backslash i}$ to denote $(X^{i-1},X_{i+1}^{n})$. The entropy of a random variable $X$ is represented as $H(X)$, while the mutual information between $X$ and $Y$ is denoted by $I(X;Y)$. We say $Q_{X}\ll P_{X}$ if $Q_{X}(x)=0$ whenever $P_{X}(x)=0$.
2 Main Results
@@
$\hat{\Psi}(X;Y\|Z)\leq\min(I(X;Y),I(X;Y|Z)).$ (11)
This follows from setting $J$ constant or $J=Z$. The main results of this section identify sufficient conditions for equality in (11). Applications will be provided in the next section.
@@
$I(T;Z)=I(X;Y|T)=I(T;J|X,Y)=I(T;J|X,Y,Z)=0.$
We have
$I(X;Y)-I(X;Y|J)$ $=I(J;X)-I(J;X|Y)$
$\leq I(J;T,X)-I(J;X|Y)$
$\leq I(J;T,X)+I(J;T,Y)-I(J;X,Y)$
=$\overset{(a)}{=}I(J;T,X)+I(J;T,Y)-I(J;X,Y,T)$
$=I(J;T,X)-I(J;X|T,Y)$
$\overset{(b)}{\leq}I(J;T)$
$\leq I(J,Z;T)$
$\overset{(c)}{\leq}I(J;T|Z)$
$\leq I(J;T,X,Y|Z)$
=$\overset{(d)}{=}I(J;X,Y|Z)$
where $(a)$ follows from $I(T;J|X,Y)=0$, $(b)$ follows from $I(X;Y|T)=0$, $(c)$ follows from $I(T;Z)=0$ and $(d)$ follows from $I(T;J|X,Y,Z)=0$. ∎
@@
$\hat{\Psi}(X;Y\|Z)=I(X;Y)$ is equivalent with the inequality
$I(X;Y|J)+I(X,Y;J|Z)\geq I(X;Y),\qquad\forall P_{J|X,Y,Z}.$
@@
$I(J;X,Y,Z)\geq\frac{1}{2}I(J;X)+\frac{1}{2}I(J;Y)+\frac{1}{2}I(J;Z),\qquad\forall P_{J|X,Y,Z}.$
This condition holds if and only if $(1/2,1/2,1/2)$ belongs to the hypercontractivity ribbon for random variables $(X,Y,Z)$ (remember that the hypercontractivity ribbon for random variables $X_{1},X_{2},\cdots,X_{k}$ is defined as the set of $\lambda_{i}$ such that for any $P_{u|x_{[k]}}$ we have $I(X_{[k]};U)\geq\sum_{i}\lambda_{i}I(X_{i};U)$).
@@
Theorem 1 readily extends to quantum systems: take quantum systems $X,Y,Z,J$ with some arbitrary joint density $\rho_{X,Y,Z,J}$. The inequality
@@
holds if there is an extended state $\rho_{X,Y,Z,J,T}$ (whose reduced density on $X,Y,Z,J$ is the given state $\rho_{X,Y,Z,J}$) such that
$I(T;Z)=I(X;Y|T)=I(T;J|X,Y)=I(T;J|X,Y,Z)=0.$
The following lemma provides an equivalent condition for the assumption of Theorem 1 for the special case of $Z$ being a function of $(X,Y)$.
@@
Consider the special case of $Z=f(X,Y)$ for some arbitrary function $f$. For a given joint distribution $P_{X,Y}$, there exists a random variable $T$ such that $I(Z;T|X,Y)=I(Z;T)=I(X;Y|T)=0$ if and only if $P_{X,Y}$ lies in the convex hull of the set
$\mathcal{A}(P_{Z})=\left\{Q_{X,Y}:I_{Q}(X;Y)=0,~{}Q_{Z}=P_{Z}\right\}.$ (12)
Proof of this lemma can be found in Appendix B.
@@ 2.1 $Z$ is a function of $(X,Y)$
In this subsection, we assume that $Z$ is a function of $X$ and $Y$. The following lemma is immediate by the choice of $U=X,Y$ and $V=Z$ in (8) which would reduce the expression to that in (9).
@@ 2.1.1 $X$ and $Y$ are binary
Let $X\oplus Y$ denote the XOR of $X$ and $Y$, and $X\wedge Y$ denote the AND of $X$ and $Y$. We consider the three functions $X+Y,X\oplus Y$ and $X\wedge Y$ as follows:
@@
  • If $Cov(X,Y)\leq 0$, then $\Psi(X;Y\|Z_{1})=\Psi(X;Y\|Z_{3})=I(X;Y)$.
  • If $Cov(X,Y)>0$, then $\Psi(X;Y\|Z_{1})=\Psi(X;Y\|Z_{3})=0.$
Proof of this theorem can be found in Appendix B.
@@ 2.1.2 Ternary-Binary
In this section, we demonstrate the applicability of Theorem 1 beyond the binary case by extending the alphabet of $Y$ to $\mathcal{Y}=\{0,1,2\}$. We give a condition for $\Psi(X;Y\|Z)=I(X;Y)$ when $Z=(X+Y)~{}mod~{}2$.
@@
Assume that $\mathcal{X}=\{0,1\},\mathcal{Y}=\{0,1,2\}$ and $\mathcal{Z}=\{0,1\}$, and
$Z=(X+Y)\mod 2=\begin{cases}0&(X,Y)\in\{(0,0),(1,1),(0,2)\}\\ 1&(X,Y)\in\{(0,1),(1,0),(1,2)\}\\ \end{cases}$ (13)
@@
Note that when $P_{Y}(2)=0$ (i.e., $Y$ is binary), $Z$ in (13), will be the XOR of binary variables $X$ and $Y$. Therefore, Theorem 3 generalizes the first part of Theorem 2.
2.1.3 Arbitary Functions
This part considers random variables $X,Y$ and $Z$ with arbitrary alphabet sizes. We aim to determine all functions $f:\mathcal{X}\times\mathcal{Y}\mapsto\mathcal{Z}$ for which the equation $\Psi(X;Y|Z)=I(X;Y)$ holds for every joint distribution $P_{X,Y}$.
@@ 2.1.3 Arbitary Functions
for some $P_{X,Y,Z}$. Then, we claim that if $X^{\prime}=f_{1}(X)$, $Y^{\prime}=f_{2}(Y)$ and $Z^{\prime}=f_{3}(Z)$ for some functions $f_{1},f_{2},f_{3}$ such that $H(Z^{\prime}|X^{\prime},Y^{\prime})=0$, then, we also have
@@
$\hat{\Psi}(X;Y\|Z)-\hat{\Psi}(X^{\prime};Y^{\prime}|Z^{\prime})\leq I(XY;Z^{\prime}|Z)+I(Y^{\prime}Z^{\prime};X|X^{\prime})+I(X^{\prime}Z^{\prime};Y|Y^{\prime})+I(X;Y|X^{\prime}Y^{\prime}Z^{\prime}).$ (14)
@@
$I(X;Y)-\hat{\Psi}(X,Y\|Z)\geq I(X^{\prime};Y^{\prime})-\hat{\Psi}(X^{\prime},Y^{\prime}\|Z^{\prime}).$ (15)
If we further have $\hat{\Psi}(X,Y\|Z)=I(X;Y)$, we deduce that
$\hat{\Psi}(X^{\prime},Y^{\prime}\|Z^{\prime})\geq I(X^{\prime};Y^{\prime}).$ (16)
@@
$\hat{\Psi}(X^{\prime},Y^{\prime}\|Z^{\prime})\leq I(X^{\prime};Y^{\prime}).$ (17)
Thus, $\hat{\Psi}(X^{\prime},Y^{\prime}\|Z^{\prime})=I(X^{\prime};Y^{\prime})$.
@@
3.1 Theorem 5, a Conjecture and Lower bounds on $S(X;Y\|Z)$
It is not known whether there exists some $P_{X,Y,Z}$ such that $S(X;Y\|Z)\neq\hat{\Psi}(X;Y\|Z)$. If $S(X;Y\|Z)=\hat{\Psi}(X;Y\|Z)$ holds for all $P_{X,Y,Z}$, we must have the following for any $P_{XYZX^{\prime}Y^{\prime}Z^{\prime}}$:
$S(X;Y\|Z)-S(X^{\prime};Y^{\prime}\|Z^{\prime})\leq I(XY;Z^{\prime}|Z)+I(Y^{\prime}Z^{\prime};X|X^{\prime})+I(X^{\prime}Z^{\prime};Y|Y^{\prime})+I(X;Y|X^{\prime}Y^{\prime}Z^{\prime}).$ (18)
@@
$S(X;Y\|Z)-S(X;Y\|Z^{\prime})\leq$ $S_{\text{ow}}(XY\rightarrow Z^{\prime}\|Z)\leq I(XY;Z^{\prime}|Z)$ (19)
@@
For any $P_{XYZX^{\prime}Y^{\prime}Z^{\prime}}$ we have
$S(X;Y\|Z)-S(X^{\prime};Y^{\prime}\|Z^{\prime})\leq S_{\text{ow}}(XY\rightarrow Z^{\prime}\|Z)+I(Y^{\prime}Z^{\prime};X|X^{\prime})+I(X^{\prime}Z^{\prime};Y|Y^{\prime})+I(X;Y|X^{\prime}Y^{\prime}Z^{\prime}).$
This conjecture generalizes (6) (it reduces to (6) by setting $X=X^{\prime}$ and $Y=Y^{\prime}$). If one can refute the conjecture, it will serve as evidence of $S(X;Y\|Z)\neq\hat{\Psi}(X;Y\|Z)$ for general distributions. If one can show the conjecture, one deduce the lower bound:
$S(X^{\prime};Y^{\prime}|Z^{\prime})\geq S(X;Y\|Z)-S_{\text{ow}}(XY\rightarrow Z^{\prime}|Z)-I(Y^{\prime}Z^{\prime};X|X^{\prime})-I(X^{\prime}Z^{\prime};Y|Y^{\prime})-I(X;Y|X^{\prime}Y^{\prime}Z^{\prime}).$
@@ 4 Is the upper bound $I(X;Y)$ tight for the XOR case?
The above inequality might potentially lead to better upper bounds because $\Psi(X^{n};Y^{n}\|Z^{n})$ allows for minimizing over all $P_{J|X^{n},Y^{n},Z^{n}}$ which is a larger space than that of the single-letter bound. The following proposition shows that at least when $Z$ is a function of $(X,Y)$, we do not obtain better upper bounds with this approach, i.e., the proposition shows that
@@
Take some arbitraray $P_{J|X^{n},Y^{n},Z^{n}}$ and let $J_{i}=(J,X^{i-1},Y^{i-1})$. Then, we have
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{I(X^{n};Y^{n}|J)}+\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{I(X^{n}Y^{n};J|Z^{n})}$ $=\sum_{i}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{I(X_{i};Y_{i}|J_{i})+I(X^{i-1};Y_{i}|JY^{i-1})+I(Y^{i-1};X_{i}|JX^{i-1})}$
$\qquad+\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{I(X_{i}Y_{i};J_{i}|Z_{i})+I(X_{i},Y_{i};Z_{\backslash i}|J_{i}Z_{i})}$
$\geq\sum_{i}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{I(X_{i};Y_{i}|J_{i})}+\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{I(X_{i}Y_{i};J_{i}|Z_{i})}$
@@
where the last equation follows from definition of $\hat{\Psi}(X;Y\|Z)$.
For the converse direction, assume that $J_{*}$ achieves the minimum of $I(X;Y|J)+I(X,Y;J|Z)$. Suppose $J$ consists of i.i.d copies of random variale $J_{*}$,i.e, $J=J_{*}^{n}$. We have
$I(X^{n};Y^{n}|J_{*}^{n})+I(X^{n}Y^{n};J_{*}^{n}|Z^{n})=n(I(X;Y|J_{*})+I(X,Y;J_{*}|Z))=n\hat{\Psi}(X;Y\|Z).$
Thus, $\hat{\Psi}(X^{n},Y^{n}\|Z^{n})\leq n\hat{\Psi}(X,Y\|Z)$. This observation and inequality (21) complete the proof. ∎
4.1 Controling $H(Z^{n}|F_{1:r})$
@@
Take some $P_{X,Y}(x,y)$ such that $P_{X,Y}(x,y)>0$ for all $x,y$. Assume that $H(Z^{n}|F_{1:r})=0$ where $F_{1:r}=(F_{1},F_{2},\cdots,F_{r})$ is the public communication. Then,
@@
[16, 17] Given two pmfs $P_{X,Y}$ and $Q_{X,Y}$ on the alphabets $\mathcal{X}\times\mathcal{Y}$, the relation $Q_{X,Y}\preceq P_{X,Y}$ represents existence of some functions $a:\mathcal{X}\mapsto\mathbb{R}_{+}\cup\{0\}$ and $b:\mathcal{Y}\mapsto\mathbb{R}_{+}\cup\{0\}$ such that for all $(x,y)\in\mathcal{X}\times\mathcal{Y}$ we have
$Q_{X,Y}(x,y)=a(x)b(y)P_{X,Y}(x,y)(x,y).$ (22)
@@
Since $H(Z^{n}|F_{1:r}=f_{1:r})=0$, the value of $z^{n}$ must be fix when we fix $F_{1:r}=f_{1:r}$. This shows that for any $x_{1}^{n},y_{1}^{n}$ and $x_{2}^{n},y_{2}^{n}$ such that $x_{1}^{n}+y_{1}^{n}\neq x_{2}^{n}+y_{2}^{n}$, the probability of $p(x_{1}^{n},y_{1}^{n}|F_{1:r}=f_{1:r})$ and $p(x_{2}^{n},y_{2}^{n}|F_{1:r}=f_{1:r})$ cannot be positive at the same time. Thus,
@@
Since $P_{X,Y}(x,y)>0$ for all $x,y$, we must have $a(x_{1}^{n})b(y_{1}^{n})a(x_{2}^{n})b(y_{2}^{n})=0$.
@@
$a(x_{1}^{n})a(x_{2}^{n})b(y^{n})=0.$
@@
$S_{\Delta}(X;Y\|Z)$ is supremum of secret key rate corresponding to protocols satisfying

### v3

@@ 1.1.2 Main Contribution 2
  • We conjecture that $S(X;Y\|Z)\neq I(X;Y)$ in the XOR case (see Section 4). Showing this would require developing a new upper bound for the source model problem. The XOR example is particularly interesting because it presents a clean case where previously known techniques – such as reducing Eve’s information via intrinsic mutual information – fail to improve the basic $I(X;Y)$ upper bound. This demonstrates the need for entirely new ideas. Partial results are also provided in our attempt to develop a new upper bound. By adapting techniques from the modulo-two sum problem in source coding, we derive an explicit bound for the non-interactive case. However, this bound is unfortunately difficult to evaluate.
We conclude the introduction with a brief historical remark. An early upper bound on $S(X;Y\|Z)$ is the intrinsic-information bound, defined in (9). It evaluates the common information between Alice’s and Bob’s sources under Eve’s most favorable choice among all of her possible “perspectives” (i.e., post-processings of $Z$). Although this bound was originally conjectured to be tight [12], Renner and Wolf [19] constructed a sophisticated counterexample showing that it can be strictly suboptimal. Conceptually, while the intrinsic-information bound weakens Eve by allowing her to degrade her observation, the upper bound in [19] moves in the opposite direction by endowing Eve with additional information. Subsequently, by introducing the idea of a fictitious terminal and developing another intricate example, [11] established the suboptimality of the bound in [19]. In this paper we show that, for the simple binary XOR source, none of these increasingly elaborate techniques for upper bounding $S(X;Y\|Z)$ improves upon the naive bound $I(X;Y)$, which is intuitively unlikely to be tight. In contrast to prior counterexamples of growing complexity, the XOR setting offers a clean and simple example; ideas developed for this source are likely to extend to more general sources as well.
The rest of this paper is organized as follows: Section 2 evaluates the upper bound $\Psi(X;Y\|Z)$ for some settings. Section 3 discusses the gap between $S(X;Y\|Z)$ and $\Psi(X;Y\|Z)$, and provides a conjecture on $S(X;Y\|Z)$. Finally, Section 4 discusses the tightness of the upper bound in the XOR case. Additionally, in Appendix A, we compute the secret key capacity for a class of sources where $Z$ is a function of $X$ and $P_{Y|X}$ corresponds to an erasure channel.
[…]
2 Optimality of constant $J$
@@
$\hat{\Psi}(X;Y\|Z)\leq I(X;Y).$ (17)
This follows from setting $J$ constant. The main results of this section identify sufficient conditions for equality in (17).
@@
I(T;Z)=I(X;Y|T)=I(T;J|X,Y)=I(T;J|X,Y,Z)=0.I(T;Z)=I(X;Y|T)=I(T;J|X,Y)=I(T;J|X,Y,Z)=0.
The proof follows from the following identity:
$I(X;Y)-\left\{I(X;Y|J)+I(X,Y;J|Z)\right\}$
$={I(T;Z)+I(X;Y|T)+I(T;Z|X,Y)+2I(T;J|X,Y,Z)}$
$\qquad-\left\{I(X;Y|J,T)+I(X,Y;J|T,Z)\right\}$
$\qquad-I(J;T|X)-I(J;T|Y)-I(T;Z|J)-I(T;Z|X,Y,J)$
$\leq 0.$
@@
$\hat{\Psi}(X;Y\|Z)=I(X;Y)$ is equivalent to the inequality
$I(X;Y|J)+I(X,Y;J|Z)\geq I(X;Y),\quad\forall P_{J|X,Y,Z}.$
@@
$I(J;X,Y,Z)\geq\frac{1}{2}I(J;X)+\frac{1}{2}I(J;Y)+\frac{1}{2}I(J;Z),\quad\forall P_{J|X,Y,Z}.$
This condition holds if and only if $(\tfrac{1}{2},\tfrac{1}{2},\tfrac{1}{2})$ belongs to the hypercontractivity ribbon for random variables $(X,Y,Z)$. Recall that the hypercontractivity ribbon for random variables $X_{1},X_{2},\ldots,X_{k}$ is defined as the set of $\lambda_{i}$ such that for any $P_{U|X_{[k]}}$ we have
$I(X_{[k]};U)\geq\sum_{i}\lambda_{i}I(X_{i};U).$
@@
Theorem 1 readily extends to quantum systems: consider quantum systems $X,Y,Z,J$ with some arbitrary joint density matrix $\rho_{X,Y,Z,J}$. The inequality
@@
holds if there exists an extended state $\rho_{X,Y,Z,J,T}$ (whose reduced density on $X,Y,Z,J$ is the given state $\rho_{X,Y,Z,J}$) such that
I(T;Z)=I(X;Y|T)=I(T;J|X,Y)=I(T;J|X,Y,Z)=0.I(T;Z)=I(X;Y|T)=I(T;J|X,Y)=I(T;J|X,Y,Z)=0.
Thus, Theorem 1 is applicable to the quantum bound presented in [20] as well (see Remark 1).
The following lemma provides an equivalent condition for the assumption of Theorem 1 in the special case where $Z$ is a function of $(X,Y)$.
@@
Consider the special case $Z=f(X,Y)$ for some arbitrary function $f$. For a given joint distribution $P_{X,Y}$, there exists a random variable $T$ such that
$I(Z;T|X,Y)=I(Z;T)=I(X;Y|T)=0$
if and only if $P_{X,Y}$ lies in the convex hull of the set
$\mathcal{A}(P_{Z})=\left\{Q_{X,Y}:I_{Q}(X;Y)=0,\quad Q_{Z}=P_{Z}\right\}.$ (18)
The proof of this lemma can be found in Appendix B.
@@ 2.1 $Z$ is a function of $(X,Y)$
In this subsection, we focus on the scenario where $Z$ is a deterministic function of $X$ and $Y$, i.e., $Z=f(X,Y)$ for some function $f$.
An immediate consequence of this assumption is that the upper bound $\Psi(X;Y\|Z)$ in (12) coincides with the simpler bound $\hat{\Psi}(X;Y\|Z)$ in (13). This follows by choosing the auxiliary random variables $U=(X,Y)$ and $V=Z$, which reduces the max-min expression in (12) directly to the form in (13).
@@ 2.1.1 $X$ and $Y$ are binary
Let $X\oplus Y$ denote the XOR of $X$ and $Y$, and $X\wedge Y$ denote the AND of $X$ and $Y$. We consider the three functions $X+Y$, $X\oplus Y$, and $X\wedge Y$ as follows:
@@
  • If $\operatorname{Cov}(X,Y)\leq 0$, then $\Psi(X;Y\|Z_{1})=\Psi(X;Y\|Z_{3})=I(X;Y).$
  • If $\operatorname{Cov}(X,Y)>0$, then $\Psi(X;Y\|Z_{1})=\Psi(X;Y\|Z_{3})=0.$
The proof of this theorem can be found in Appendix B.
@@ 2.1.2 Ternary-Binary
In this section, we demonstrate the applicability of Theorem 1 beyond the binary case by extending the alphabet of $Y$ to $\mathcal{Y}=\{0,1,2\}$. We give a condition for $\Psi(X;Y\|Z)=I(X;Y)$ when $Z=(X+Y)\mod 2$.
@@
Assume that $\mathcal{X}=\{0,1\}$, $\mathcal{Y}=\{0,1,2\}$, and $\mathcal{Z}=\{0,1\}$, and
$Z=(X+Y)\mod 2=\begin{cases}0&(X,Y)\in\{(0,0),(1,1),(0,2)\}\\ 1&(X,Y)\in\{(0,1),(1,0),(1,2)\}\end{cases}$ (19)
@@
Note that when $P_{Y}(2)=0$ (i.e., $Y$ is binary), $Z$ in (19) will be the XOR of binary variables $X$ and $Y$. Therefore, Theorem 3 generalizes the first part of Theorem 2.
2.1.3 Arbitrary Functions
As we observed earlier, when $Z$ is the XOR of $X$ and $Y$, we have $\Psi(X;Y\|Z)=I(X;Y)$ for every source $P_{X,Y}$. Equivalently, the set $\mathcal{P}$ (defined in the introduction) restricted to $Z=X\oplus Y$ is the entire set of all joint distributions on $\mathcal{X}\times\mathcal{Y}$. Is there any function, other than XOR, with this property? This part considers random variables $X,Y,$ and $Z$ with arbitrary alphabet sizes. We aim to determine all functions $f:\mathcal{X}\times\mathcal{Y}\mapsto\mathcal{Z}$ for which the equation $\Psi(X;Y\|Z)=I(X;Y)$ holds for every joint distribution $P_{X,Y}$.
@@ 2.1.3 Arbitrary Functions
for some $P_{X,Y,Z}$. Then, we claim that if $X^{\prime}=f_{1}(X)$, $Y^{\prime}=f_{2}(Y)$ and $Z^{\prime}=f_{3}(Z)$ for some functions $f_{1},f_{2},f_{3}$ such that $H(Z^{\prime}|X^{\prime},Y^{\prime})=0$, then we also have
@@
$\hat{\Psi}(X;Y\|Z)-\hat{\Psi}(X^{\prime};Y^{\prime}\|Z^{\prime})\leq I(X,Y;Z^{\prime}|Z)+I(Y^{\prime},Z^{\prime};X|X^{\prime})+I(X^{\prime},Z^{\prime};Y|Y^{\prime})+I(X;Y|X^{\prime},Y^{\prime},Z^{\prime}).$ (20)
@@
$I(X;Y)-\hat{\Psi}(X;Y\|Z)\geq I(X^{\prime};Y^{\prime})-\hat{\Psi}(X^{\prime};Y^{\prime}\|Z^{\prime}).$ (21)
If we further have $\hat{\Psi}(X;Y\|Z)=I(X;Y)$, we deduce that
$\hat{\Psi}(X^{\prime};Y^{\prime}\|Z^{\prime})\geq I(X^{\prime};Y^{\prime}).$ (22)
@@
$\hat{\Psi}(X^{\prime};Y^{\prime}\|Z^{\prime})\leq I(X^{\prime};Y^{\prime}).$ (23)
Thus, $\hat{\Psi}(X^{\prime};Y^{\prime}\|Z^{\prime})=I(X^{\prime};Y^{\prime})$.
@@
3.1 Remarks on Theorem 5, a Conjecture and Lower bounds on $S(X;Y\|Z)$
It is not known whether there exists some $P_{X,Y,Z}$ such that $S(X;Y\|Z)\neq\hat{\Psi}(X;Y\|Z)$. If $S(X;Y\|Z)=\hat{\Psi}(X;Y\|Z)$ holds for all $P_{X,Y,Z}$, we must have the following for any $P_{X,Y,Z,X^{\prime},Y^{\prime},Z^{\prime}}$:
$S(X;Y\|Z)-S(X^{\prime};Y^{\prime}\|Z^{\prime})\leq I(X,Y;Z^{\prime}|Z)+I(Y^{\prime},Z^{\prime};X|X^{\prime})+I(X^{\prime},Z^{\prime};Y|Y^{\prime})+I(X;Y|X^{\prime},Y^{\prime},Z^{\prime}).$ (24)
@@
$S(X;Y\|Z)-S(X;Y\|Z^{\prime})\leq$ $S_{\text{ow}}(X,Y\rightarrow Z^{\prime}\|Z)\leq I(X,Y;Z^{\prime}|Z).$ (25)
@@
For any $P_{X,Y,Z,X^{\prime},Y^{\prime},Z^{\prime}}$, we have
$S(X;Y\|Z)-S(X^{\prime};Y^{\prime}\|Z^{\prime})\leq S_{\text{ow}}(X,Y\rightarrow Z^{\prime}\|Z)+I(Y^{\prime},Z^{\prime};X|X^{\prime})+I(X^{\prime},Z^{\prime};Y|Y^{\prime})+I(X;Y|X^{\prime},Y^{\prime},Z^{\prime}).$
This conjecture generalizes (10) (it reduces to (10) by setting $X=X^{\prime}$ and $Y=Y^{\prime}$). If one can refute the conjecture, it will serve as evidence of $S(X;Y\|Z)\neq\hat{\Psi}(X;Y\|Z)$ for general distributions. If one can show the conjecture, one deduces the lower bound:
$S(X^{\prime};Y^{\prime}\|Z^{\prime})\geq S(X;Y\|Z)-S_{\text{ow}}(X,Y\rightarrow Z^{\prime}\|Z)-I(Y^{\prime},Z^{\prime};X|X^{\prime})-I(X^{\prime},Z^{\prime};Y|Y^{\prime})-I(X;Y|X^{\prime},Y^{\prime},Z^{\prime}).$
@@ 4 Is the upper bound $I(X;Y)$ tight for the XOR case?
The above inequality might potentially lead to better upper bounds because $\Psi(X^{n};Y^{n}\|Z^{n})$ allows for minimizing over all $P_{J|X^{n},Y^{n},Z^{n}}$, which is a larger space than that of the single-letter bound. The following proposition shows that at least when $Z$ is a function of $(X,Y)$, we do not obtain better upper bounds with this approach, i.e., the proposition shows that
@@
Take some arbitrary $P_{J|X^{n},Y^{n},Z^{n}}$ and let $J_{i}=(J,X^{i-1},Y^{i-1})$. Then, we have
$I(X^{n};Y^{n}|J)+I(X^{n},Y^{n};J|Z^{n})$ $=\sum_{i}\bigl{(}I(X_{i};Y_{i}|J_{i})+I(X^{i-1};Y_{i}|J,Y^{i-1})+I(Y^{i-1};X_{i}|J,X^{i-1})\bigr{)}$
$\quad+\sum_{i}\bigl{(}I(X_{i},Y_{i};J_{i}|Z_{i})+I(X_{i},Y_{i};Z^{i-1},Z_{i+1}^{n}|J_{i},Z_{i})\bigr{)}$
$\geq\sum_{i}\bigl{(}I(X_{i};Y_{i}|J_{i})+I(X_{i},Y_{i};J_{i}|Z_{i})\bigr{)}$
@@
where the last inequality follows from the definition of $\hat{\Psi}(X;Y\|Z)$.
For the converse direction, assume that $J_{*}$ achieves the minimum of $I(X;Y|J)+I(X,Y;J|Z)$. Suppose $J$ consists of i.i.d. copies of the random variable $J_{*}$, i.e., $J=J_{*}^{n}$. We have
$I(X^{n};Y^{n}|J_{*}^{n})+I(X^{n}Y^{n};J_{*}^{n}|Z^{n})=n\bigl{(}I(X;Y|J_{*})+I(X,Y;J_{*}|Z)\bigr{)}=n\hat{\Psi}(X;Y\|Z).$
Thus, $\hat{\Psi}(X^{n};Y^{n}\|Z^{n})\leq n\hat{\Psi}(X;Y\|Z)$. This observation and inequality (27) complete the proof. ∎
4.1 Controlling $H(Z^{n}|F_{1:r})$
@@
Take some $P_{X,Y}(x,y)$ such that $P_{X,Y}(x,y)>0$ for all $x,y$. Assume that $H(Z^{n}|F_{1:r})=0$, where $F_{1:r}=(F_{1},F_{2},\cdots,F_{r})$ is the public communication. Then,
@@
[29, 30] Given two pmfs $P_{X,Y}$ and $Q_{X,Y}$ on the alphabets $\mathcal{X}\times\mathcal{Y}$, the relation $Q_{X,Y}\preceq P_{X,Y}$ represents the existence of some functions $a:\mathcal{X}\mapsto\mathbb{R}_{+}\cup\{0\}$ and $b:\mathcal{Y}\mapsto\mathbb{R}_{+}\cup\{0\}$ such that for all $(x,y)\in\mathcal{X}\times\mathcal{Y}$ we have
$Q_{X,Y}(x,y)=a(x)b(y)P_{X,Y}(x,y).$ (28)
@@
Since $H(Z^{n}|F_{1:r}=f_{1:r})=0$, the value of $z^{n}$ must be fixed when we fix $F_{1:r}=f_{1:r}$. This shows that for any $x_{1}^{n},y_{1}^{n}$ and $x_{2}^{n},y_{2}^{n}$ such that $x_{1}^{n}+y_{1}^{n}\neq x_{2}^{n}+y_{2}^{n}$, the probabilities $p(x_{1}^{n},y_{1}^{n}|F_{1:r}=f_{1:r})$ and $p(x_{2}^{n},y_{2}^{n}|F_{1:r}=f_{1:r})$ cannot be positive at the same time. Thus,
@@
Since $P_{X,Y}(x,y)>0$ for all $x,y$, we must have
$a(x_{1}^{n})b(y_{1}^{n})a(x_{2}^{n})b(y_{2}^{n})=0.$
@@
$a(x_{1}^{n})a(x_{2}^{n}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}b^{2}(y^{n})}=0.$
@@
$S_{\Delta}(X;Y\|Z)$ is the supremum of secret key rates corresponding to protocols satisfying